OpenStudy (anonymous):

r^4 + 2r^2 + r = 0 ... How to solve this ???

OpenStudy (anonymous):

factor out r

OpenStudy (anonymous):

r (r^3 +2r +1 )

OpenStudy (anonymous):

for the second term use factor theorm

OpenStudy (anonymous):

note, if you put r=-1 in the second bracket then it is zeroed out so (r+1) is a factor

OpenStudy (anonymous):

therefore (r^3 +2r +1 ) = (r+1)(r^2 +ar ) let r=1 4 = 2(1+a) 1+a = 2 a= 1

OpenStudy (anonymous):

umm if i put r = -1 in the second bracket the answer doesn't come out to be zero :S

OpenStudy (anonymous):

ohh yeh, made mistake there

OpenStudy (anonymous):

umm so what am i supposed to do , i can't think of any number at which it becomes zero

OpenStudy (anonymous):

I think I know something that might work

OpenStudy (anonymous):

let f(r) = r^4 +2r^2 +r f ' (r) = 4r^3 +4r = 4r ( r^2 +1 ) which as only one zero at r=0 , and it can be shown by the first derivative test that this is a minimum

OpenStudy (anonymous):

so , when r=0 when have a minimum turning point , and since r=0 is a zero of the function that means the function oly has one zero ( at r=0 ) so r=0 is the only real solution